OpenStudy (anonymous):

1. Create your own third degree polynomial that when divided by x + 2 has a remainder of –4. 2. Create your own division of polynomials problem. Demonstrate how this problem would be solved using both long division and synthetic division.

OpenStudy (anonymous):

what is the degree if you multiply x^2*(x+2) ?

OpenStudy (anonymous):

\[\huge x^2(x+2) = \]

OpenStudy (anonymous):

just multiply it out...

OpenStudy (anonymous):

not quite... \[\large x^2(x+2)=x^3+2x \]... agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that's a third degree polynomial.. and what is the remainder if we take \(\large x^3+2x \) divided by \(\large x+2 \) ???

OpenStudy (anonymous):

remember we just multiplied it out to get that third degree polynomial...

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

confused...................

OpenStudy (anonymous):

if I ask you to multipl 2 times 3, what's the product?

OpenStudy (anonymous):

**multiply...

OpenStudy (anonymous):

6

OpenStudy (anonymous):

now... what's the remainder if I ask you 6 divided by 3 ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

that's what i was asking in when we were working with the polynomial x^3 + 2x... the remainder when \(\large x^3+2x \) divided by \(\large x+2 \) is zero....

OpenStudy (anonymous):

still with me?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

good... but we want a remainder of -4.... so what do you think we need to do?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):